Exercise 9.20 Constant Part A A compact disc (CD) stores music in a coded pattern of tiny pits 107 m deep. The pits are arranged in a track that spirals outward toward the rim of the disc the inner and outer radil of this spiral are 25.0 mm and 58.0 mm, respectively. As the disc spins inside a CD player, the track is scanned at a constant linear speed of 1.25 m/s. What is the angular speed of the CD when scanning the innermost part of the track? rad/s Request Answer Part B What is the angular speed of the CD when scanning the outermost part of the track? rad/s Part C playing time of a CD is 74.0 min. What would be the length of the track on such a maximu out in a
media%2Febd%2Febd98b98-8894-4762-a00c-30
Answer #1

A) use, w = v/r

= 1.25/(25*10^-3)

= 50 rad/s

B) w = v/r

= 1.25/(58*10^-3)

= 21.6 rad/s

C) L = v*t

= 1.25*74*60

= 5550 m

= 5.55 km

D) alfa = (w2 - w1)/t

= (21.6 - 50)/(74*60)

= -0.00640 rad/s^2